# Simple Controller Design to Achieve Iso-Damping Robustness: Non-Iterative Data-Driven Approach Based on Fractional-Order Reference Model

### v2

@@ II-A The flat-phase property and iso-damping
where $\angle L\left(s\right)$ represents the argument of $L\left(s\right)$ [18]. If $L\left(s\right)$ in Fig. 1 has the flat-phase property, the closed-loop system shown in Fig. 1 has a special robustness: the overshoot amount of the output $y$ to the setpoint reference input $r$ is invariant under the variation of $K$, namely, the iso-damping [4]. The iso-damping property is practically useful since the plant gain variation often occurs in many control systems.
@@
A typical choice of $L\left(s\right)$ realizing the flat-phase property is the BITF $L_{Bode}\left(s\right)$ defined as
@@
where $\gamma\in\mathbb{R}$. Note that the BITF is an FO transfer function since the FO integrator/differentiator becomes $s^{\alpha}\;(\alpha\in\mathbb{R})$ in the Laplace domain ($\alpha<0$ for integrator, $\alpha>0$ for differentiator). This fact clearly shows that the FO control framework is effective to realize the iso-damping property, owing to its high design flexibility. The design policy for the BITF has been discussed in [18]. $\blacksquare$
Figure 1: Closed-loop system.
@@ II-B BIBO stability of proper rational discrete-time IO SISO LTI system
Let $\left\{g_{k}\right\}_{k=0}^{\infty}$ be the impulse response sequence of $G\left(z\right)$ and let $\left\{u_{k}\right\}_{k=0}^{\infty}$ be an arbitrary bounded input sequence, i.e., there exists $0<\alpha_{u}<\infty$ such that $\left\lvert u_{k}\right\rvert<\alpha_{u}\;(k=0,1,2,\ldots)$. Then, $G\left(z\right)$ is said to be BIBO stable if there exists $0<\alpha_{y}<\infty$ satisfying $\left\lvert g_{k}\ast u_{k}\right\rvert<\alpha_{y}\;(k=0,1,2,\ldots)$. $\blacksquare$
@@
    All poles of $G\left(z\right)$ exist inside the unit circle. $\blacksquare$
(Proof.) See Section 4.3.2 in [49]. ∎
@@ II-C Overview of the controller design problem in this study
Here, the input data $u_{[0:N]}^{D}=\left\{u_{k}^{D}\right\}_{k=0}^{N}$ and the output data $y_{[0:N]}^{D}=\left\{y_{k}^{D}\right\}_{k=0}^{N}$ are collected by using the reference input $r_{[0:N]}^{D}=\left\{r_{k}^{D}\right\}_{k=0}^{N}\;\left(r_{0}^{D}\neq 0\right)$ and some initial controller $C\left(z;\theta_{0}\right)$ under the zero initial condition. In this study, we address the problem of finding $\theta$ such that $T\left(z;\theta\right)$ satisfies the designer-specified design specification including the iso-damping property, where $P\left(z\right)$ is unknown. Throughout this study, the input/output data is assumed to be noiseless. Under noisy conditions, the total variation denoising [50] is a practical choice to deal with the measurement noise.
@@
III The proposed approach
This section describes the controller tuning problem described in Section II-C so as to achieve the iso-damping robust control.
@@ III-A Robust controller design via MR-D3C framework
Fig. 3 sketches the proposed controller design approach based on the MR-D3C framework. Let $L_{flat}\left(s\right)$ be the continuous-time FO SISO LTI transfer function satisfying (2), and let $L_{flat}\left(s\right)$ be such that $L_{flat}\left(z\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(L_{flat}\left(s\right)\right)$. In Fig. 3, $M_{ref}\left(z\right)=L_{flat}\left(z\right)\left\{1+L_{flat}\left(z\right)\right\}^{-1}$ is the FC-based reference model specifying the desired closed-loop characteristics. Note that $M_{ref}\left(z\right)$ has the iso-damping robustness owing to the specification of $L_{flat}\left(z\right)$. The designer a priori determines the specific choice of $L_{flat}\left(s\right)$. The meaning of $P\left(z\right)$, $C\left(z;\theta\right)$, $r$, $u$, and $y$ are the same as those in Fig 2. The following assumption can be easily satisfied by suitably designing $L_{flat}\left(z\right)$.
@@
$L_{flat}\left(z\right)$ is proper and rational. $M_{flat}\left(z\right)$ has all poles inside the unit circle, implying $M_{flat}\left(z\right)$ is BIBO stable. $\blacksquare$
@@
Find the optimal parameter $\theta^{\ast}\in\mathbb{R}^{n}$ of the controller on the basis of $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$, such that $T\left(z;\theta^{\ast}\right)r_{[0:N]}^{D}$ is as close to $M_{ref}\left(z\right)r_{[0:N]}^{D}$ as possible in the mean square error sense. $\blacksquare$
@@
Since $L_{flat}\left(z\right)$ is constructed based on the FO transfer function, it is preferable that the class of the controller is also chosen as the FC-based controller such as $C\left(z;\theta\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(C_{FOPID}\left(s;\theta\right)\right)$, where $C_{FOPID}\left(s;\theta\right)$ is the standard FO-PID controller with the tunable parameter $\theta$. This fact is demonstrated in Example 1 in Section IV. $\blacksquare$
Figure 3: MR-D3C approach to design the controller so as to achieve the iso-damping robustness.
@@ III-B Fictitious reference signal
The proposed approach addresses Problem 1 based on the fictitious reference signal [37, 51]. The fictitious reference signal $\tilde{r}_{[0:N]}\left(\theta\right)=\left\{\tilde{r}_{k}\left(\theta\right)\right\}_{k=0}^{N}$ is defined as
@@
For any $\theta$ such that $1+P\left(z\right)C\left(z;\theta\right)\neq 0$, $T\left(z;\theta\right)\tilde{r}_{[0:N]}\left(\theta\right)=y_{[0:N]}^{D}$. $\blacksquare$
(Proof) See Eq. (8) in [37]. ∎
@@
One study has proposed a MR-D3C technique based on the fictitious reference signal: FRIT [37]. However, the controller provided by FRIT may destabilize the closed-loop system as pointed out in [50, 52]. To overcome this difficulty, several studies have examined to modify FRIT for IO control [50, 52]. $\blacksquare$
III-C The proposed controller design approach
@@
[…]
$J\left(\theta\right)=\left(\left\lVert R^{D}\left(\tilde{R}\left(\theta\right)\right)^{-1}\mathrm{vec}\left(y_{[0:N]}^{D}\right)\right.\right.\$ (6)
where $R^{D}=\mathrm{Tpl}\circ\mathrm{vec}\left(r_{[0:N]}^{D}\right)$ and $\tilde{R}\left(\theta\right)=\mathrm{Tpl}\circ\mathrm{vec}\left(\tilde{r}_{[0:N]}\left(\theta\right)\right)$. $\blacksquare$ (Proof) Let $t\left(\theta\right)=\left\{t_{k}\left(\theta\right)\right\}_{k=0}^{\infty}$ be the impulse response sequence of $T\left(z;\theta\right)$. Due to Lemma 1,
$\mathrm{vec}\left(y^{D}\right)=\tilde{R}\left(\theta\right)\mathrm{vec}\left(t_{[0:N]}\left(\theta\right)\right)$ (7)
where $t_{[0:N]}\left(\theta\right)\coloneqq\left\{t_{k}\left(\theta\right)\right\}_{k=0}^{N}$. Since $\tilde{r}_{0}\left(\theta\right)\neq 0$, the impulse response of $T\left(z;\theta\right)$, which is the closed-loop transfer function from the reference input to the output when the controller parameter is $\theta$, from time 00 to $N$ is restored as
@@
Therefore, $y_{[0:N]}\left(\theta\right)\coloneqq\left\{y_{k}\left(\theta\right)\right\}_{k=0}^{N}$, i.e., the output of $T\left(z;\theta\right)$ due to $r_{[0:N]}^{D}$, can be estimated as
@@
$\mathrm{vec}\left(M^{ref}\left(z\right)r_{[0:N]}^{D}\right)=R^{D}\mathrm{vec}\left(m_{[0:N]}^{ref}\right)$ (10)
@@
Equation (1) shows that the loss function $J\left(\theta\right)$ in the minimization problem (5) explicitly reflects the impulse response of $T\left(z;\theta\right)$. If the value of $J\left(\theta\right)$ is reasonably small, we can consider that $T\left(z;\theta\right)$ is BIBO stable. Contrarily, the controller making the value of $J\left(\theta\right)$ excessively large should destabilize the closed-loop system. That is, we can evaluate the BIBO stability of the resultant closed-loop system from a bounded reference input to a bounded output before implementing the controller. Therefore, iso-IDFRIT can design a reliable and practically feasible controller. $\blacksquare$
@@ III-D Stability analysis
The optimization problem in (5) is successfully solved for the asymptotic case. That is, $J\left(\theta\right)\to\alpha$ for $N\to\infty$, where $\alpha\in\mathbb{R}_{+}$ is a finite value. $\blacksquare$
The following lemma is the instrumental for the stability analysis:
@@
Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI transfer function. The impulse response sequence of $G\left(z\right)$ is denoted as $g=\left\{g_{k}\right\}_{k=0}^{\infty}$. Then, $G\left(z\right)$ is BIBO stable if and only if $\left\lVert g\right\rVert_{2}<\infty$. $\blacksquare$
(Proof) Let $(A,B,C,D)$ be the the minimal state-space realization of $G\left(z\right)$, i.e., $G(z)=C\left(zI-A\right)^{-1}B+D$ ($I$ is the identity matrix with the compatible dimension). Then, $g_{0}=D$ and $g_{k}=CA^{k-1}B$ for $k\geq 1$. Note that $\Lambda\left(A\right)$ coincides the poles of $G\left(z\right)$ including multiplicity due to the minimality of $(A,B,C,D)$ (Theorem 12.9.16 in [53]). This implies that $G\left(z\right)$ is BIBO stable if and only if $\rho\left(A\right)<1$ due to (C2) in Proposition 1.
Assume that $G\left(z\right)$ is BIBO stable. Then, due to (C1) in Proposition 1, $\left(\left\lVert g\right\rVert_{2}\right)^{2}<\left\lVert g\right\rVert_{\infty}\left\lVert g\right\rVert_{1}<\infty$. In contrast, suppose that $G\left(z\right)$ is not BIBO stable, i.e., $\rho\left(A\right)\geq 1$. Here, for $k\to\infty$, $\left\lvert g_{k}\right\rvert^{2}=\left\lvert CA^{k-1}B\right\rvert^{2}\to 0$ if and only if $A^{k-1}\to 0$. However, $A^{k-1}\to 0$ if and only if $\rho\left(A\right)<1$ (Theorem 5.6.12 in [54]), proving that $\left\lVert g\right\rVert_{2}$ is not finite. ∎
@@
Under the Assumptions 1 and 2, $T\left(z;\theta^{\ast}\right)$ is BIBO stable. $\blacksquare$
(Proof) Define
@@
Then,
. (12)
That is, Here, $J\left(\theta^{\ast}\right)=\left(\left\lVert\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right\rVert_{2}\right)^{2}=\left(\left\lVert\mathrm{vec}\left(\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}\right)^{2}$ due to (8) and (LABEL:Eq_Vec-epsilon). Therefore,
@@
In Theorem 2, Assumption 2 is a sufficient condition for proving the BIBO stability of $T\left(z;\theta^{\ast}\right)$; the BIBO stability may be achieved even if Assumption 2 does not hold. Therefore, the stability result in Theorem 2 suffers from some conservatism. Moreover, Theorem 2 is the asymptotic result. Analysis in the finite sample case and the derivation of the necessary and sufficient condition of the BIBO stability are left as future work. Nevertheless, the proposed loss function $J\left(\theta\right)$ in (6) explicitly reflects the pole information of $T\left(z;\theta^{\ast}\right)$ via its $\left(N+1\right)$-length impulse response $t_{[0:N]}\left(\theta\right)$. This feature allows us to infer that the resultant controller $C\left(z;\theta^{\ast}\right)$ stabilizes the closed-loop system from the reference input to the output in the BIBO sense, improving the reliability of the tuning result. $\blacksquare$
Remark 7.
We can interpret the loss function $J\left(\theta\right)$ from the viewpoint of the frequency domain. Specifically, since $J\left(\theta\right)=\sum_{n=1}^{N}\left\lvert\left(t_{k}\left(\theta\right)-m_{k}^{ref}\right)\ast r_{k}^{D}\right\rvert^{2}$,
where $\Phi_{rD}\left(e^{j\omega}\right)$ is the spectral density of the reference input $r_{[0:\infty]}^{D}$ used in the initial experiment. Therefore, the minimization of $J\left(\theta\right)$ corresponds to the minimization of the mismatch of the frequency responses $T\left(e^{j\omega};\theta\right)$ and $M_{ref}\left(e^{j\omega}\right)$ with the frequency weighting function $\Phi_{rD}\left(e^{j\omega}\right)$. $\blacksquare$

### v3

@@ II-A Flat-phase property and iso-damping
where $\angle L\left(s\right)$ represents the argument of $L\left(s\right)$ [4]. If $L\left(s\right)$ in Fig. 1 has the flat-phase property, the closed-loop system shown in Fig. 1 has a special robustness: the overshoot amount of the output $y$ to the setpoint reference input $r$ is invariant under the variation of $K$, namely, the iso-damping [4]. The iso-damping property is practically useful, as the plant gain variation often occurs in many control systems.
Figure 1: Closed-loop system.
@@
A typical choice of $L\left(s\right)$ realizing the flat-phase property is the BITF $L_{Bode}\left(s\right)$:
@@
where $\gamma\in\mathbb{R}$. Note that the BITF is an FO transfer function because the FO integrator/differentiator becomes $s^{\alpha}\;(\alpha\in\mathbb{R})$ in the Laplace domain ($\alpha<0$ for integrator, $\alpha>0$ for differentiator). This fact clearly shows that the FO control framework is effective to realize the iso-damping property, owing to its high design flexibility. The design policy for the BITF has been discussed in [20]. Further details of the Laplace transforms of FO derivatives and integrals can be found, for example, in Section 2.8 of [53].
@@ II-B BIBO stability of proper rational discrete-time IO SISO LTI system
Let $\left\{g_{k}\right\}_{k=0}^{\infty}$ be the impulse response sequence of $G\left(z\right)$ and let $\left\{u_{k}\right\}_{k=0}^{\infty}$ be an arbitrary bounded input sequence, i.e., there exists $0<\alpha_{u}<\infty$ such that $\left\lvert u_{k}\right\rvert<\alpha_{u}\;(k=0,1,2,\ldots)$. Then, $G\left(z\right)$ is said to be BIBO stable if there exists $0<\alpha_{y}<\infty$ satisfying $\left\lvert g_{k}\ast u_{k}\right\rvert<\alpha_{y}\;(k=0,1,2,\ldots)$.
@@
    All poles of $G\left(z\right)$ exist inside the unit circle.
Proof: See Section 4.3.2 in [52]. ∎
@@ II-C Overview of the controller design problem in this study
Here, the input data $u_{[0:N]}^{D}=\left\{u_{k}^{D}\right\}_{k=0}^{N}$ and the output data $y_{[0:N]}^{D}=\left\{y_{k}^{D}\right\}_{k=0}^{N}$ of the controlled plant $P\left(z\right)$ are collected through an open-loop or closed-loop experiment. In this study, we address the problem of finding $\theta$ such that $T\left(z;\theta\right)$ satisfies the designer-specified design specification including the iso-damping property, where $P\left(z\right)$ is unknown. Throughout this study, the input/output data is assumed to be noiseless. Under noisy conditions, the total variation denoising [54] is a practical choice to deal with the measurement noise.
@@
III Proposed approach
This section presents the controller tuning problem described in Section II-C in order to achieve the iso-damping robust control.
@@ III-A Robust controller design via MR-D3C framework
Fig. 3 sketches the proposed controller design approach based on the MR-D3C framework. Let $L_{flat}\left(s\right)$ be the continuous-time FO SISO LTI transfer function satisfying (2), and $L_{flat}\left(z\right)\coloneq\mathrm{c2d}\circ\mathrm{f2i}\left(L_{flat}\left(s\right)\right)$. In Fig. 3, $M_{ref}\left(z\right)=L_{flat}\left(z\right)\left\{1+L_{flat}\left(z\right)\right\}^{-1}$ is the FC-based reference model specifying the desired closed-loop characteristics. Note that $M_{ref}\left(z\right)$ has the iso-damping robustness owing to the specification of $L_{flat}\left(z\right)$. The designer a priori determines the specific choice of $L_{flat}\left(s\right)$. The meaning of $P\left(z\right)$, $C\left(z;\theta\right)$, $r$, $u$, and $y$ are the same as those in Fig 2. The following assumption can be easily satisfied by suitably designing $L_{flat}\left(z\right)$.
@@
$L_{flat}\left(z\right)$ is proper and rational. $M_{ref}\left(z\right)$ has all poles inside the unit circle, implying $M_{ref}\left(z\right)$ is BIBO stable.
@@
Find the optimal parameter $\theta^{\ast}\in\mathbb{R}^{n}$ of the controller on the basis of $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$, such that $T\left(z;\theta^{\ast}\right)r_{[0:N]}^{D}$ is as close to $M_{ref}\left(z\right)r_{[0:N]}^{D}$ as possible in the mean square error sense. Here, $r_{[0:N]}^{D}=\left\{r_{k}^{D}\right\}_{k=0}^{N}\,\left(r_{0}^{D}\neq 0\right)$ is the reference input specified by the designer.
@@
As $L_{flat}\left(z\right)$ is constructed based on the FO transfer function, it is preferable that the class of the controller is also chosen as the FC-based controller such as $C\left(z;\theta\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(C_{FOPID}\left(s;\theta\right)\right)$, where $C_{FOPID}\left(s;\theta\right)$ is the standard FO-PID controller with the tunable parameter $\theta$. This fact is demonstrated in Example 1 in Section IV.
Figure 3: MR-D3C approach to design a controller to achieve the iso-damping robustness.
@@ III-B Fictitious reference signal
The proposed approach addresses Problem 1 using the fictitious reference signal [39, 55]. The fictitious reference signal $\tilde{r}_{[0:N]}\left(\theta\right)=\left\{\tilde{r}_{k}\left(\theta\right)\right\}_{k=0}^{N}$ is defined as
@@
For any $\theta$ such that $1+P\left(z\right)C\left(z;\theta\right)\neq 0$, $T\left(z;\theta\right)\tilde{r}_{[0:N]}\left(\theta\right)=y_{[0:N]}^{D}$.
Proof: The statement follows from straightforward calculations [39]. In particular,
$T\left(z;\theta\right)\tilde{r}_{[0:N]}\left(\theta\right)$ (5)
$=\frac{P\left(z\right)C\left(z;\theta\right)}{1+P\left(z\right)C\left(z;\theta\right)}\left\{\left(C\left(z;\theta\right)\right)^{-1}u_{[0:N]}^{D}+y_{[0:N]}^{D}\right\}$ (6)
$=\frac{1}{1+P\left(z\right)C\left(z;\theta\right)}P\left(z\right)u_{[0:N]}^{D}+\frac{P\left(z\right)C\left(z;\theta\right)}{1+P\left(z\right)C\left(z;\theta\right)}y_{[0:N]}^{D}$ (7)
$=\frac{1}{1+P\left(z\right)C\left(z;\theta\right)}\left\{y_{[0:N]}^{D}+P\left(z\right)C\left(z;\theta\right)y_{[0:N]}^{D}\right\}$ (8)
$=y_{[0:N]}^{D}$ (9)
which is the desired outcome. ∎
@@
One study has proposed a MR-D3C technique based on the fictitious reference signal: FRIT [39]. However, the controller provided by FRIT may destabilize the closed-loop system as pointed out in [54, 56]. To overcome this difficulty, several studies have examined to modify FRIT for IO control [54, 56].
III-C Proposed controller design approach
@@
$\theta^{\ast}=\underset{\theta}{\arg\min}\>J\left(\theta\right)$ (10)
where $R^{D}=\mathrm{Tpl}\circ\mathrm{vec}\left(r_{[0:N]}^{D}\right)$ and $\tilde{R}\left(\theta\right)=\mathrm{Tpl}\circ\mathrm{vec}\left(\tilde{r}_{[0:N]}\left(\theta\right)\right)$. Proof: Let $t\left(\theta\right)=\left\{t_{k}\left(\theta\right)\right\}_{k=0}^{\infty}$ be the impulse response sequence of $T\left(z;\theta\right)$. Due to Lemma 1,
$\mathrm{vec}\left(y_{[0:N]}^{D}\right)=\tilde{R}\left(\theta\right)\mathrm{vec}\left(t_{[0:N]}\left(\theta\right)\right)$ (12)
where $t_{[0:N]}\left(\theta\right)\coloneqq\left\{t_{k}\left(\theta\right)\right\}_{k=0}^{N}$. Since $\tilde{r}_{0}\left(\theta\right)\neq 0$, the impulse response of $T\left(z;\theta\right)$, which is the closed-loop transfer function from the reference input to the output when the controller parameter is $\theta$, from time 0 to $N$ is restored as
@@
Here, let $y_{[0:N]}\left(\theta\right)=\left\{y_{k}\left(\theta\right)\right\}_{k=0}^{N}$ denote the output of $T\left(z;\theta\right)$ due to $r_{[0:N]}^{D}$. Using (13), $y_{[0:N]}\left(\theta\right)$ can be estimated as
@@
$\mathrm{vec}\left(M_{ref}\left(z\right)r_{[0:N]}^{D}\right)=R^{D}\mathrm{vec}\left(m_{[0:N]}^{ref}\right)$ (15)
@@
Equation (III-C) shows that the loss function $J\left(\theta\right)$ in the minimization problem (10) explicitly reflects the impulse response of $T\left(z;\theta\right)$. If the value of $J\left(\theta\right)$ is reasonably small, we can consider that $T\left(z;\theta\right)$ is BIBO stable. Conversely, the controller making the value of $J\left(\theta\right)$ excessively large should destabilize the closed-loop system. That is, we can evaluate the BIBO stability of the resultant closed-loop system from a bounded reference input to a bounded output before implementing the controller. Therefore, iso-IDFRIT can design a reliable and practically feasible controller.
@@ III-D Stability analysis
The optimization problem (10) is successfully solved for the asymptotic case. That is, $J\left(\theta\right)\to\alpha$ for $N\to\infty$, where $\alpha\in\mathbb{R}_{+}$ is a finite value.
The following lemma is instrumental for the stability analysis:
@@
Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI transfer function. The impulse response sequence of $G\left(z\right)$ is denoted as $g=\left\{g_{k}\right\}_{k=0}^{\infty}$. Then, $G\left(z\right)$ is BIBO stable if and only if $\left\lVert g\right\rVert_{2}<\infty$.
Proof: Let $(A,B,C,D)$ be the the minimal state-space realization of $G\left(z\right)$, i.e., $G(z)=C\left(zI-A\right)^{-1}B+D$ (II is the identity matrix with the compatible dimension). Then, $g_{0}=D$ and $g_{k}=CA^{k-1}B$ for $k\geq 1$. Note that $\Lambda\left(A\right)$ coincides the poles of $G\left(z\right)$ including multiplicity due to the minimality of $(A,B,C,D)$ (Theorem 12.9.16 in [57]). This implies that $G\left(z\right)$ is BIBO stable if and only if $\rho\left(A\right)<1$ due to Proposition 1.
Assume that $G\left(z\right)$ is BIBO stable. Then, due to Proposition 1, $\left(\left\lVert g\right\rVert_{2}\right)^{2}<\left\lVert g\right\rVert_{\infty}\left\lVert g\right\rVert_{1}<\infty$. In contrast, suppose that $G\left(z\right)$ is not BIBO stable, i.e., $\rho\left(A\right)\geq 1$. Here, for $k\to\infty$, $\left\lvert g_{k}\right\rvert^{2}=\left\lvert CA^{k-1}B\right\rvert^{2}\to 0$ if and only if $A^{k-1}\to 0$. However, $A^{k-1}\to 0$ if and only if $\rho\left(A\right)<1$ (Theorem 5.6.12 in [58]), proving that $\left\lVert g\right\rVert_{2}$ is not finite. ∎
@@
Under the Assumptions 1 and 2, $T\left(z;\theta^{\ast}\right)$ is BIBO stable.
Proof: Let $\varepsilon_{[0:N]}\left(\theta^{\ast}\right)$ denote the error between the output of the closed-loop system $T\left(z;\theta^{\ast}\right)$ driven by $r_{[0:N]}^{D}$ and the desired closed-loop response due to $r_{[0:N]}^{D}$. That is,
@@
Then, That is, Here,
$J\left(\theta^{\ast}\right)=\left\lVert\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right\rVert_{2}^{2}=\left\lVert\mathrm{vec}\left(\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}^{2}$ (19)
due to (13) and (LABEL:Eq_Vec-epsilon). Therefore,
@@
In Theorem 2, Assumption 2 is a sufficient condition for proving the BIBO stability of $T\left(z;\theta^{\ast}\right)$; the BIBO stability may be achieved even if Assumption 2 does not hold. Moreover, Theorem 2 is the asymptotic result. Analysis in the finite sample case and the derivation of the necessary and sufficient condition of the BIBO stability are left as future work. Nevertheless, the proposed loss function $J\left(\theta\right)$ in (LABEL:Eq_Loss_iso-IDFRIT) explicitly reflects the pole information of $T\left(z;\theta^{\ast}\right)$ via its $\left(N+1\right)$-length impulse response $t_{[0:N]}\left(\theta\right)$. This feature allows us to infer that the resultant controller $C\left(z;\theta^{\ast}\right)$ stabilizes the closed-loop system from the reference input to the output in the BIBO sense, improving the reliability of the tuning result.